OpenStudy (anonymous):

Is there anyone who can help me with this problem... (advanced algebra) (x^2-25)/x^5+x^4+18x^3+18x^2-175x-175 I'm trying to find the zeros. I have to do this first... Any help?

jimthompson5910 (jim_thompson5910):

Hint: if P/Q = 0, then P/Q = 0 P = 0*Q P = 0 So if P/Q = 0, then P = 0 The value of Q, it turns out, is completely irrelevant. As long as Q is not zero, then everything should be fine.

OpenStudy (anonymous):

Um.... I'm sorry but I'm even more confused by that answer.

jimthompson5910 (jim_thompson5910):

P and Q can be any number or expression. So in this case P = x^2 - 25 and Q = x^5+x^4+18x^3+18x^2-175x-175 So (x^2-25)/(x^5+x^4+18x^3+18x^2-175x-175) becomes P/Q

OpenStudy (anonymous):

ah perhaps I should have mentioned that I am supposed to be solving this by long division and therein lies my problem. I am terrible at long division. Especially with problems like that.

jimthompson5910 (jim_thompson5910):

There's no need to use long division though since only the numerator matters.

jimthompson5910 (jim_thompson5910):

Basically, if (x^2-25)/(x^5+x^4+18x^3+18x^2-175x-175) = 0, then x^2 - 25 = 0

jimthompson5910 (jim_thompson5910):

Does that make sense?

jimthompson5910 (jim_thompson5910):

If not, then think of something like this If (x+1)/(x+2) = 0, then... (x+1)/(x+2) = 0 x+1 = 0*(x+2) ... multiply both sides by x+2 x+1 = 0x+0*2 x + 1 = 0 + 0 x + 1 = 0 So if (x+1)/(x+2) = 0, then x + 1 = 0 This shows us that the denominator plays no role whatsoever in finding the zeros.

OpenStudy (anonymous):

\[x^{2} -25\div x^{5}+x^{4}+18^{3}+18^{2}-175x-175\] That's not what I'm looking for though. This is just a piece of the big problem I have to solve. What I need to do is solve this equation, and use the answer to find the zeros of the original.

jimthompson5910 (jim_thompson5910):

That's the problem. What you just wrote isn't an equation.

jimthompson5910 (jim_thompson5910):

Also, I'm assuming that x^5+x^4+18x^3+18x^2-175x-175 is all in the denominator correct?

OpenStudy (anonymous):

That is the dividend. the x^2-25 is the divisor. The point is that I have to legitimately divide the equation in order to finish the overall problem.

jimthompson5910 (jim_thompson5910):

Oh so it's really something like